OpenStudy (anonymous):

integral of sqrt(1+x^2)/x^2

OpenStudy (anonymous):

What do you think you can do?

OpenStudy (anonymous):

trig substitution probably.

OpenStudy (anonymous):

Yes. The identity of 1 + tan^2(x) applies. How do you substitute x = tan(x)?

OpenStudy (anonymous):

sub x = tanx, then dx = \[\sec ^{2}x\] Then: int of \[\sqrt(1+x^2)/x^2\] = \[(\sqrt(\tan^2x + 1)(\sec^2x)/\tan^2x)\] = \[\int\limits_{}^{} \sec^3x/\tan^2x. \] Not sure what to do now.

OpenStudy (anonymous):

Um, sqrt(tan^2(x)+ 1) = sec(x), not sec^2(x). Anyway, sec(x) = 1/cos x, tan x = sin x / cos x. Substitute.

OpenStudy (anonymous):

Nevermind, I see your logic.

OpenStudy (anonymous):

Or do I?

OpenStudy (anonymous):

\[\sec^2x\] is from the dx. dx = sec^2x

OpenStudy (anonymous):

Yeah, thought so. Continue.

OpenStudy (anonymous):

\[\int\limits_{}^{}(1+\tan^2x(secx))/\tan^2x = \int\limits_{}^{} secx/tanx + secxtan^2x/\tan^2x = \int\limits_{}^{} secx/tanx + secx = \int\limits_{}^{} sinx/\cos^2x + secx = 1/cosx + \ln \left| secx + tanx \left| \right| \right|\]

OpenStudy (anonymous):

how do you make the screen bigger